OpenStudy (anonymous):

3/4+x=7/8 Solve? I dun get it D:

OpenStudy (anonymous):

Wait wait wait would x = 4/4 or would it be 1 7/8? or yeah I'm dumb in math.

OpenStudy (anonymous):

Why would it be 7-6? :O

OpenStudy (anonymous):

because 8 in 8 is 1 then 1*7=7 - 4in 8 is 2 2*3=6 so you have 7-6 :P

OpenStudy (anonymous):

Oh okay so would it be 1/4 if I wanted to simplify it or what?

OpenStudy (anonymous):

no it's just 1/8

OpenStudy (anonymous):

1/8 is end

OpenStudy (anonymous):

Alrighty thanks